A 10 kg weight is raised to a geight of 2 m with an acceleration of 1.5 metre per secobd square .Compute the work done if g=10 metre per second square .

Grade:

1 Answers

Mahima Kanawat
5 years ago
Dear student 
Work done W = F. d
Here F is N + ma= mg ( lift is moving upwards  so, N and accelerating force are in same direction) 
N = m (g - a)
N = 10 (10 - 1.5 )
N= 85
and d= height=2m
Therefore W = 85*2 = 170 N
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT